OpenStudy (anonymous):

Expand the binomial:

OpenStudy (anonymous):

\[(\sqrt{x}-\sqrt{y})^{5}\]

OpenStudy (perl):

i will expand it

OpenStudy (perl):

( x^1/2 - y^1/2 ) ^5 = we use pascal triangle coefficients 1 5 10 10 5 1

OpenStudy (perl):

I am going to let a = x^1/2, and b = y^1/2 , then we can go back and substitute ( x^1/2 - y^1/2 ) ^5 = a^5 + 5a^4*b + 10a^3*b^2 + 10a^2* b^3 + 5a*b^4 + b^5

OpenStudy (anonymous):

\[(a-b)^{5} = a ^{5}-5a ^{4}b+10a ^{3}b ^{2}-10a ^{2}b ^{3}+5ab ^{4}-b ^{5}\] now replace: \[a = \sqrt{x} \] \[b = \sqrt{y}\]

OpenStudy (perl):

you did the same thing i did, wow. what a waste of time

OpenStudy (perl):

and you used pretty print too.

OpenStudy (anonymous):

@perl you forgot your negatives

OpenStudy (anonymous):

that doesnt help me.....

OpenStudy (anonymous):

@laurenbae what will help you?

OpenStudy (anonymous):

5 i x^(3/2) sqrt(y) e^(i pi floor(-(arg(y))/(2 pi))) see here http://www.wolframalpha.com/input/?i=%28x%E2%88%9A%E2%88%92y%E2%88%9A%295&dataset=

OpenStudy (anonymous):

but that's not one of the choices it gives me >.<

OpenStudy (mimi_x3):

This might be an easier approach than the pascals triangle.. \[(a+b)^{n} = \binom{n}{k}*a^{n-k}*b^{k}\] \(( \sqrt{x} - \sqrt{y})^5\) \[ \sum_{k=0}^{5}\binom{5}{k} * \left(\sqrt{x}\right)^{5-k} *\left(-\sqrt{y}\right)^{k}\]

OpenStudy (mimi_x3):

Are you able to do it?

OpenStudy (anonymous):

no which is why im asking for help. >.<

OpenStudy (mimi_x3):

where are you having problems?